OpenStudy (pixiedust1):

Evaluate -b+sqrtb^2-4ac for the given values of a, b, and c. Write the result as a complex number.

OpenStudy (pixiedust1):

\[-b+\sqrt{b^2-4ac}\] a=1 b=-6 c=13

OpenStudy (pixiedust1):

I got \[6+2i \sqrt{22}\] and I'm nearly positive that it's correct, just wanted to see if you could double-check it since double negatives always throw me :)

OpenStudy (anonymous):

i got \[-6+4i\]

OpenStudy (zenmo):

\[\frac{ -(-6)\pm \sqrt{(-6)^2-4(1)(13)} }{ 2(1) } = \frac{ 6 \pm \sqrt{36-52} }{ 2 }\] \[\frac{ 6\pm \sqrt{-16} }{ 2 } = \frac{ 6 \pm 4i }{ 2 }= 3 \pm 2i = 3 + 2i, 3 - 2i\]

OpenStudy (anonymous):

my bad 6 plus and minus 4i, unless it's asking you to solve the whole quadratic equation with 2a in the denominator, in which case zenmo is correct. If it is just askin -b+...., then the answer is 6+4i and 6-4i,

OpenStudy (pixiedust1):

So my answer is correct?

OpenStudy (pixiedust1):

@jhonyy9 Hey mind looking at this question and checking to see if my answer is correct?

OpenStudy (pixiedust1):

@Kainui Hey, would you mind looking at this question and checking to see if my answer is correct?

jhonyy9 (jhonyy9):

a=1 b=-6 c=13 -b + sqrt(b^2 -4ac) -(-6) +sqrt(36-4*13*1 6 +sqrt(36-52) 6+sqrt(-16) 6+4i you have got this ?

OpenStudy (pixiedust1):

I see, so it's 6+4i and not 6+2i sqrt 22?

jhonyy9 (jhonyy9):

6+4i is correct sure

OpenStudy (pixiedust1):

Thank you Jhonyy!

jhonyy9 (jhonyy9):

was my pleasure and anytime you like or i m here online bye bye

jhonyy9 (jhonyy9):

and good luck